Ed has 95 cents in dimes and nickels. The total number of coins is one more than twice the number of dimes. How many dimes does she have?
I'm confused on this because it doesn't doesn't tell me a total number so I'm
Not sure what to go off of
The only equation I could think of was
.10d + .05n = .95
Not the other one we should have if someone could help
Me find the other equation and find the amount of coins.
Please and thank you !

Found 2 solutions by Alan3354, ankor@dixie-net.com:
Answer by Alan3354(69443)

Ed has 95 cents in dimes and nickels. The total number of coins is one more than twice the number of dimes. How many dimes does she have?
=======================
5n + 10d = 95
n+d = 2d+1

Answer by ankor@dixie-net.com(22740)

Ed has 95 cents in dimes and nickels. The total number of coins is one more than twice the number of dimes. How many dimes does she have?
I'm confused on this because it doesn't doesn't tell me a total number so I'm
Not sure what to go off of
The only equation I could think of was
.10d + .05n = .95
:
Write an equation for the statement
"the total number of coins is one more than twice the number of dimes."
d + n = 2d + 1
n = 2d - d + 1
n = d + 1
now, in your first equation you can substitute (d+1) for n
.10d + .05(d+1) = .95
.10d + .05d + .05 = .95
.15d = .95 - .05
.15d = .90
d = .90/.15
d = 6 dimes
I'll let you find the no. of nickels, check your solutions in the 1st equation